OpenStudy (anonymous):

The sum of the interior angles of a quadrilateral is: 540° 720° 360° 180°

OpenStudy (johnweldon1993):

Use the general rule of \[\large (n - 2) \times 180^{\circ}\] so for a quadrilateral (4 sides) we have \[\large (4 - 2) \times 180^{\circ}\] \[\large 2 \times 180^{\circ}\] \[\large 360^{\circ}\]

OpenStudy (anonymous):

the answer is 360 there is an equation that looks like this: (n-2)*180 and n stands for the number of angles